solve the following eqn for x where [x] and {x} denote greatest integer and fractional part


 


2x + 3[x] - 4{-x}=4


plz give detaild soln

Grade:12th Pass

1 Answers

Sher Mohammad IIT Delhi
9 years ago
2x+3[x]-4+4{x}=4
2x+3x+{x}=8
5[x]+6{x}=8
{x}=(8-5[x])/6
that means 0<=(8-5[x])/6<1
0<=8-5[x]<6
-8<=-5[x]<-2
2<5[x]<=8
2/5<[x]<=8/5
[x]=0,1
0<=x<2

Sher Mohammad
faculty askiitians